OpenStudy (anonymous):

find the area of a square with a side length of 8 inches. A) A = 64√3 sq. in. B) A = 65 sq. in. C) A = 64 cu. sq. in D) None of the above

OpenStudy (btaylor):

A=s^2 A=8^2=64 square inches.